# Accessing Semi-Addressable Self Assembly with Efficient Structure Enumeration

### v1

@@ Abstract
Using our structure enumeration algorithm, we have shown that both assembly economy and quality can be improved by venturing from the fully-addressable into the semi-addressable assembly regime. While other authors have found semi-addressable designs that prevent off-target structures [38, 15], for example by taking advantage of symmetries or repeated structural motifs, our results show that off-target structures do not have to be avoided to still achieve high equilibrium yield. This observation has profound implications for exploiting semi-addressability for multifarious self assembly, where a single set of building blocks assembles into multiple target structures simultaneously [14, 19, 15], or for any use case where building blocks need to be reused. Furthermore, we are able to achieve these high-yield designs while adhearing to the experimentally relevant constraints of equal binding energies and stochiometric concentrations. Additionally optimizing over binding energies and monomer concentrations should lead to further increases in target yield [14], but is not necessary in the structures we study.
This result is enabled by our encoding of structures as vertex-colored directed graphs, which we call their anatomy. This anatomy allows us to efficiently define the functions $f^{+}(s)$ and $f^{-}(s)$, which are the necessary ingredients for implementing reverse search. Together, this leads to an enumeration algorithm with minimal memory requirements that is so computationally efficient that we were able to perform inverse design through a quasi-brute-force approach (Fig. 4). Combining our fast enumeration and yield calculations with more efficient methods for generating candidate assembly systems, such as the techniques of Bohlin et al. [15], would allow semi-addressable inverse design of larger and more complex structures.
@@
Our work is accompanied by an open-source, user-orientated implementation of our enumeration algorithm [34] that can be used for any 2D or 3D system with orientationally locking interactions. This implementation could be further scaled up through large-scale parallelization, made possible by reverse search [29], or extended beyond orientationally-locking interactions by generalizing the raising and lowering functions, as discussed in the SI. Finally, structure enumeration can be critical for addressing a wide range of topics in self assembly beyond the scope of this work, including questions of kinetics [39, 40, 41, 42] or hierarchy [9, 16, 17]. For example, the structure graph (Fig. 2c) could be exploited for Markov-based kinetic calculations, or hierarchical assembly schemes could be implemented by reusing high-yield output structures as building blocks for a second round of self assembly. We envision our enumeration algorithm to be a flexible tool for many such applications.

### v2

@@ Abstract
These results show that both assembly economy and quality can be improved by decreasing the number of species, thus moving from the fully-addressable into the semi-addressable assembly regime. While other authors have found semi-addressable designs that prevent off-target structures, for example by exploiting symmetry [12, 13, 29] or through other computational approaches [42, 21], our results show that off-target structures do not have to be avoided to still achieve high equilibrium yield. This observation has profound implications for exploiting semi-addressability for multifarious self assembly, where a single set of building blocks assembles into multiple target structures simultaneously [24, 25, 21], or for any use case where building blocks need to be reused. Furthermore, we are able to achieve these high-yield designs while adhering to the experimentally relevant constraints of equal binding energies and stoichiometric concentrations. Additionally optimizing over binding energies and particle concentrations should lead to further increases in target yield [20], but is not necessary for the structures we study.
This result is enabled by our enumeration algorithm, which is so computationally efficient that we were able to perform inverse design through a quasi-brute-force approach (Fig. 3). Combining our fast enumeration and yield calculations with more efficient methods for generating candidate assembly systems, such as the techniques of Bohlin et al. [21], should allow semi-addressable inverse design of larger and more complex structures. However, care must be taken when considering structures that self-close on length scales large compared to particle size, such as tubules [11, 27, 28], shells [10], or other complex surfaces [13]. In these cases, even small amounts of bond flexibility can lead to the formation of strained off-target structures that violate the assumption of perfectly rigid bonds, and suppressing such off-target structures generally requires sacrificing some assembly economy [11, 27, 28, 13]. Our algorithm could be modified in future work to include bond flexibility (see SI), which would extend our method to this more general class of systems.
@@
Finally, we note that structure enumeration can be critical for addressing a wide range of topics in self assembly beyond the scope of this work. For example, the structure network (Fig. 2c) could be exploited for kinetic calculations [43, 44, 45, 46], or hierarchical assembly schemes [16, 22, 23] could be implemented by reusing high-yield output structures as building blocks for a second round of self assembly. We envision our enumeration algorithm, which is freely available online [47], to be a flexible tool for many such applications.